# Fact-checking information generated by a large language model can decrease news discernment

### v3

@@ Discussion
While our experimental design allows us to assess the causal effects of LLM fact-checking information on the discernment of true and false news, it is important to exercise caution when generalizing these results to different contexts. First, we use a specific version of ChatGPT to generate fact-checking information; the findings may not apply to other AI models. Second, design choices intended to emulate a realistic social media environment — such as including headline sources and lede text — may contribute to people’s assessments, although these effects should be equal for all experimental conditions. Third, the survey setting of our experiment may not fully capture the complexities of real-world information consumption and sharing behaviors. However, previous research has shown a correlation between self-reported willingness to share news in online surveys and actual sharing behavior on social media platforms53. Finally, while our study presents real news headlines that replicate a common social media design, the results may not generalize beyond our relatively small selection of political news headlines. Nevertheless, the pretest conducted on these articles ensured they are balanced with respect to dimensions known to be important to believing and sharing news (see Methods).
Despite these limitations, our study provides valuable insights into the complex interplay between humans and AI in the context of automated fact checking. ChatGPT performs well at debunking false headlines while it mostly reports being unsure about true headlines, consistent with previous research49, 5, 34. Although an ideal model would, of course, be largely correct, the LLM’s ability to clearly communicate uncertainty (without specific instruction to do so) may be considered a desirable quality of an automated fact-checking system54. Furthermore, ChatGPT is a general-purpose model; given the rapid development of AI technology, performance in fact checking is likely to improve as language models are developed explicitly for this task.
@@
The average belief and sharing discernment of participants were not affected by viewing LLM fact-checking information, whether or not such information was optional. These results are surprising given previous work suggests that LLMs can persuade humans on controversial topics55, 56. However, we found that AI-generated fact checks can affect belief in and intent to share news headlines, and this effect is contingent upon the accuracy of the AI’s responses relative to the veracity of the headlines. Consistent with literature showing that AI may be perceived as objective38, 39, participants tended to believe true headlines less when the LLM incorrectly labeled them as false. Conversely, when the LLM expressed uncertainty about the veracity of false headlines, participants were more inclined to believe them. We also found that participants were more likely to share false headlines in this scenario, although the effect was less significant. Confirmation bias may partially explain these results57, but further research is needed to understand the underlying mechanisms. Participants also demonstrated an increased willingness to share true headlines that were correctly identified by the LLM. While most of our results indicate undesirable behavioral changes, this outcome is encouraging, as it aligns with efforts to improve the acceptance of reliable information58.
@@
The behavior of participants in the optional condition revealed a strong selection bias. When participants were given the choice to view LLM fact-checking information, those who chose to do so were significantly more likely to share both true and false news, but only more likely to believe false news. These results suggest that individuals may have already formed their opinion about a headline before accessing the fact-checking information. For example, they might wish to confirm what they believe to be true or see if the AI is wrong. Irrespective of the motivation, they subsequently disregard the fact check. This pattern is particularly evident with respect to false headlines, for which ChatGPT provides highly accurate information. Despite being presented with helpful information indicating that these headlines were false, participants were still much more likely to report believing or being willing to share that content. Further interaction analyses suggest that individual attitudes towards AI, as well as partisan congruence with headlines, are related to this behavior. Although our study design cannot reveal the exact mechanism behind the outcomes of the optional condition, the findings suggest that this misinformation intervention design is unlikely to be helpful. An interesting extension of this experiment would be to test an intervention in which users are not aware that fact-checking information is AI-generated.
@@
We present these results in the context of concerns raised by experts about the potential for AI to contribute to the digital misinformation problem59, 60, 61, 62. These concerns are well-founded; malicious AI-powered bots are virtually undetectable on social media63 and even the developers of ChatGPT report that their technology is likely to be weaponized by malicious actors64, 60. To make matters worse, recent research suggests that state-of-the-art LLMs can persuade individuals on polarized topics56, 65, providing an incentive for their use in political information campaigns60.
@@
As artificial intelligence becomes increasingly integrated into our information environment, it is crucial that we thoroughly understand the risks as well as the opportunities it creates. Our results reveal that the dynamics of human-AI interaction make the application of LLMs to fact checking potentially harmful despite its accuracy.
@@ Methods
We utilized Qualtric’s quota-matching system to ensure that our sample was representative of the United States population with respect to gender, age, race, education, and partisanship. We utilized 2020 U.S. Census66 and Pew Research67 data as references for our quota criteria, which Qualtrics guaranteed with a $\pm$5accuracy. Further details can be found in Supplementary methods. After sampling, data for 1,548 participants were collected. In the belief group, the control, optional, and forced conditions had 241, 261, and 247 participants, respectively. In the sharing group, the control, optional, and forced conditions had 267, 263, and 269 participants, respectively. We conducted $\chi^{2}$ tests to compare the distributions of sharing and accuracy groups (control vs. optional vs. forced) across the above dimensions and found no significant differences. All subjects confirmed their consent to participate in this study, which was approved by Indiana University’s IRB (protocol 1307012383).
@@
We utilize 40 headlines with a balance of perceived partisanship, impact, familiarity, sensationalism, and the likelihood of being shared and believed. Half were true and half false. Each story included a headline, a lede sentence (if present), and an image. All headline stimuli are included in our preregistration48. Further details can be found in Supplementary methods.
@@
Fact-checking information was generated by submitting to ChatGPT the prompt “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” This prompt was designed to capture a realistic scenario in which someone uses an AI chatbot to fact-check a headline to which they were exposed. All fact checks are included in our preregistration48. Further details can be found in Supplementary methods. To quantify and account for ChatGPT’s fact-checking accuracy, the first three authors independently labeled the fact-checking information as either “True,” “Unsure,” or “False.” Final annotations were based on the majority labels (Krippendorff’s $\alpha=$ 0.79).
@@
All participants began by completing a brief survey, followed by exposure to their respective experimental conditions, followed by another brief survey and debriefing. Regardless of the condition, all participants saw the same headlines in random order. These stimuli were presented simultaneously with fact-checking information or questions about viewing fact checks (depending on experimental condition) along with questions regarding individual belief and sharing intention. Participants who failed an attention check were excluded from the study. Further details can be found in Supplementary methods.
@@ Acknowledgements
The authors would like to acknowledge Lisa Fazio, David Rand, Stephan Lewandowsky, Adam Berinsky, Mark Susmann, and Gordon Pennycook who kindly shared the news headline stimuli developed for the Mercury Project “Building a better toolkit (for fighting misinformation)”68. This research was supported in part by Knight Foundation and Volkswagen Foundation.
@@ 1.1 Sampling details
Our goal was to obtain .95 power to detect a small effect size of .1 at the standard .05 error probability with two-by-three-level between-subject manipulations ([Accuracy vs. Sharing groups] $\times$ [Control, Forced, Option conditions]). Power analysis by the G*Power69 software suggested a minimum number of 44 subjects per condition ($N=264$), but we aimed for a larger target sample size of $N=1{,}500$ (250 participants per condition) to increase the precision of our measurements.
In our final sample, females comprised 52.45% of the sample, males 47.42%, and other genders 0.13%. Age segments were 65+ (20.35%), 55-64 (16.67%), 45-54 (17.76%), 35-44 (16.73%), 25-34 (17.25%), and 18-24 (11.24%). Race percentages were: White (60.40%), Hispanic or Latino/a (17.31%), Black or African American (13.37%), Asian (5.04%), and Other (3.88%). Slightly more than half of the sample (54.46%) had less than a college education, while 45.54% had a college degree. With respect to party identification, 48.20% identified as Democrat or Democrat-leaning, 42.76% as Republican or Republican-leaning, 6.65% as Independent, and 0.26% as Other, while 1.81% reported having no preference, and 0.32% reported not knowing their affiliation.
@@ 1.2 Stimuli curation
We selected the news stories used as stimuli from a project aimed at comparing misinformation interventions 68. Specifically, 40 headlines were selected from a set of 317 political news stories using a pretest approach70, 71, 72 to balance the selected headlines in terms of perceived partisanship, impact, familiarity, sensationalism, and the likelihood of being shared and believed.
@@
The claims were presented in a digital format resembling popular social media platforms, commonly known as the “Facebook format”50.
1.3 Fact check generation
@@
A new ChatGPT session was created on the publicly available OpenAI website (chat.openai.com), where the headline text was inserted into a prompt asking, “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” The source of an article (e.g., “nytimes.com”) was excluded. If an article’s lede sentence was shown in the stimulus image, it was also included in the prompt, separated by a colon. The prompt for each headline was provided to ChatGPT only once, and the response was saved as a screenshot. All headlines were generated on January 25, 2023, between 12:30–8:00pm Eastern Standard Time. According to the release notes73, the language model utilized by ChatGPT at that time was a version of GPT-3.5 that has since been updated and is no longer available.
1.4 Attention check
@@
We incorporated an attention-check question that involved a headline created by the researchers stating that the color of the sky is yellow. Prior to viewing any headlines, participants were informed about this specific headline and instructed to later answer “Yes” when asked if they believed the headline or were willing to share it, depending on their respective experimental conditions. To minimize the distinction between the attention check and the regular experimental stimuli, this question was formatted in the same manner as all other headlines. This attention check headline was then presented randomly within the 40 stimuli headlines. Participants who answered this question incorrectly were excluded from the final data sample provided by Qualtrics.
@@ 2.2 Attitude towards AI
Questions were answered with a seven-point Likert scale ranging from “strongly disagree” to “strongly agree.” Items 1 and 3 were reverse coded such that higher values on all items indicated greater trust in artificial intelligence. For our regression analyses (see Section Regression analyses for details), each participant’s ATAI is calculated as the mean value of their responses to this battery.
@@ 2.3 Headline congruence
A headline is considered “congruent” with a participant’s partisan perspective if it is typically considered to be favorable towards the political party that they are affiliated with. Headlines are either pro-Democrat or pro-Republican, based on the pretest described in the main text. Thus, a congruent headline for a Democrat (Republican) would be one that is pro-Democrat (pro-Republican). Conversely, an incongruent headline for a Democrat (Republican) would be one that is pro-Republican (pro-Democrat).
We estimated participant partisanship by asking participants the following question: “Generally speaking do you think of yourself as a Republican, a Democrat, an Independent, or what?” Possible answers were “Democrat,” “Republican,” “Independent,” “No Preference,” “Don’t know,” and “Other” (with a text box to fill if this option is selected). If a participant selected “Independent” as their answer to this question they were then asked, “Do you think of yourself as closer to the Republican or Democratic Party?” Possible answers were “Republican Party,” “Democratic Party,” “Don’t know,” and “Neither.” We consider participants who answered “Democrat” for the first question or “Democratic Party” for the second question as Democrats. We consider as Republicans those who answered “Republican” for the first question or “Republican Party” for the second question. In other words, Independents who lean towards Democrats (Republicans) were recoded as Democrats (Republicans) in our analysis.

### v4

@@ Discussion
While our experimental design allows us to assess the causal effects of LLM fact-checking information on the discernment of true and false headlines, it is important to exercise caution when generalizing these results to different contexts. First, we use a specific version of ChatGPT to generate fact-checking information with a single prompt; these results may not apply to other AI models or prompting approaches. Although our prompt aimed to reflect naturalistic usage, its realism is uncertain due to the lack of prior research on how people use LLMs for fact-checking in real-world settings. Second, design choices intended to emulate a realistic social media environment—such as including headline sources and lede text—may contribute to people’s assessments, although these effects should be equal for all experimental conditions. Third, the survey setting of our experiment may not fully capture the complexities of real-world information consumption and sharing behaviors. However, previous research has shown a correlation between self-reported willingness to share news in online surveys and actual sharing behavior on social media platforms (54). Finally, while our study presents real headlines that replicate a common social media design, the results may not generalize beyond our relatively small selection of political news. Nevertheless, the pretest conducted on these headlines ensured they are balanced with respect to dimensions known to be important to believing and sharing news (see Methods).
Despite these limitations, our study provides valuable insights into the complex interplay between humans and AI in the context of automated fact checking. ChatGPT performs well at identifying false headlines while it mostly reports being unsure about true headlines, consistent with previous research (32, 33, 34). Since we tested a limited number of headlines, it is difficult to determine why the model judges false headlines more accurately. The model may be more likely to have seen false headline stimuli as their publication dates were less recent than those of true headlines. This highlights a key limitation of large-scale automated fact-checking systems that we refer to as the “breaking news problem”: developing news stories often discuss novel events the model has never been exposed to, making it difficult for AI to assess them accurately. To this end, a promising future research direction is to augment LLMs with trusted data—e.g., via real-time search (35)—to improve their performance on new and evolving information (55).
@@
While the average belief and sharing discernment of participants was positively affected by viewing human fact checks, this was not the case for viewing LLM fact-checking information, whether or not such information was optional. These results are surprising, considering previous research suggests that LLMs can persuade humans on controversial topics (56). However, we found that AI-generated fact checks can affect belief in and intent to share news headlines, contingent upon the accuracy of the AI’s responses relative to the veracity of the headlines. Consistent with literature showing that AI may be perceived as objective (40, 41), participants tended to believe true headlines less when the LLM incorrectly labeled them as false. Furthermore, participants demonstrated an increased willingness to share true headlines that were correctly identified by the LLM. The latter outcome is encouraging, as it supports efforts to enhance the acceptance of reliable information (57). Since trusted content is far more abundant than misinformation, future research should investigate how the volume of different types of content interacts with model accuracy to impact overall information quality.
When the LLM expressed uncertainty about the veracity of false headlines, participants were more inclined to believe and share them. This contradicts research suggesting that uncertain fact checks can be perceived as false (58), and that expressions of uncertainty from an LLM can increase task accuracy (59). While expressing uncertainty has been considered a desirable quality in automated fact-checking systems (60), our results illustrate that unsure fact checks can lead to adverse outcomes. Given the impact of the format of fact checks (61, 62), this conflicting evidence highlights an important question for future research: which formats and styles of AI-generated fact checks are most effective, and which prompting techniques can reliably create them?
@@
The behavior of participants in the optional condition revealed a strong selection bias. When participants were given the choice to view LLM fact-checking information, those who chose to do so were significantly more likely to share both true and false news. Furthermore, those who viewed this information were less likely to believe true news misjudged as false and more likely to believe false news. These results suggest that individuals may have already formed their opinion about a headline before accessing the fact-checking information. For example, they might wish to confirm what they believe to be true or see if the AI is wrong. Of course, many factors may influence how one seeks and processes fact-checking information, including how well-informed (63) and confident they are (64). Regardless, some participants subsequently disregard these fact checks. This pattern is particularly evident with respect to false headlines, for which ChatGPT provides highly accurate information. Despite being presented with helpful information indicating that these headlines were false, participants were still much more likely to report believing or being willing to share that content. Further interaction analyses suggest that individual attitudes towards AI, as well as partisan congruence with headlines, are related to this behavior. Although our study design cannot reveal the exact mechanism behind the outcomes of the optional condition, the findings suggest that this misinformation intervention design is unlikely to be helpful.
Future work could explore the effect of telling people whether a fact check comes from a human or AI. Similar questions have been investigated in the context of generic conversations (65), health prevention (66), advertising (67), and written content (68). In these scenarios, disclosing the AI-generated source tends to lead to a negative perception of the content and a preference for human-generated content. A dedicated investigation on the effect of fact-checking source disclosure will be required.
@@
We present these results in the context of concerns raised by experts about the potential for AI to contribute to the digital misinformation problem (69, 70, 71, 72). These concerns are well-founded; malicious AI-powered bots are virtually undetectable on social media (73) and even the developers of ChatGPT report that their technology is likely to be weaponized by malicious actors (74, 70). To make matters worse, recent research indicates that state-of-the-art LLMs can persuade individuals on polarized topics (56, 75) and create persuasive propaganda (76), providing an incentive for their use in political information campaigns (70).
@@
While the use of LLM-powered fact-checking to combat these concerns is enticing, our results reveal that the dynamics of human-AI interaction make this application potentially harmful, despite its accuracy. This should not discourage us from exploring the potential of this technology to help us mitigate challenging problems. Instead, as artificial intelligence becomes more deeply integrated into our information environment, it is crucial to fully understand both the risks and opportunities it presents.
@@ Methods
We utilized Qualtric’s quota-matching system to ensure that our sample matched the United States population with respect to gender, age, race, education, and partisanship. We utilized 2020 U.S. Census (77) and Pew Research (78) data as references for our quota criteria, which Qualtrics guaranteed with a $\pm$5accuracy. We conducted $\chi^{2}$ tests to compare the distributions across the above dimensions for each experimental condition (control vs. LLM-optional vs. LLM-forced vs. human fact check), for both belief and sharing groups. We find one significant difference: in the belief group only, participants in the human fact check condition were more educated (i.e., held degrees) than those in the LLM-optional condition. Our analyses do not make comparisons between these two groups and our main results are confirmed by regression analyses that account for this and other factors. Further details can be found in the Supplementary Information. After sampling, data for 2,159 participants were collected. In the belief group, the control, LLM-optional, LLM-forced, and human fact check conditions had 241, 261, 247, and 300 participants, respectively. In the sharing group, the control, LLM-optional, LLM-forced, and human fact check conditions had 267, 263, 269, and 311 participants, respectively. The drop-out rate was low (between 1%–6%) across experimental conditions (see Supplementary Information). All subjects confirmed their consent to participate in this study, which was approved by Indiana University’s IRB (protocol 1307012383).
The data for the control, LLM-optional, and LLM-forced conditions were collected in March 2023. At the request of reviewers, data for the human fact check conditions were gathered later, from March to June 2024. All data were collected following the same protocols. Participants were randomly assigned to one of the conditions at their respective times of collection.
@@
We utilize 40 real news headlines that are related to US politics, balanced in terms of partisanship, believability, and the likelihood of being shared. These headlines were generated for another study (79). Half were true and half false. Each story included a headline, a lede sentence (if present), and an image. All headline stimuli are included in our preregistration (50). Further details can be found in the Supplementary Information.
@@
Fact-checking information was generated by submitting to ChatGPT the prompt “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” This prompt was designed to capture a realistic scenario in which someone uses an AI chatbot to fact-check a headline to which they were exposed. All fact checks are included in our preregistration (50). To quantify and account for ChatGPT’s fact-checking accuracy, the first three authors independently labeled the fact-checking information as either “True,” “Unsure,” or “False.” Final annotations were based on the majority labels (Krippendorff’s $\alpha=$ 0.79). Further details can be found in the Supplementary Information.
Human fact checks
Each human fact check begins with a clear statement about the truthfulness of the claim, such as “The above claim is True” or “The above claim is False.” Following this, the fact check addresses the publisher’s reputation: if the headline is true, it mentions that the publisher is trustworthy; if false, it highlights the publisher’s unreliability. Brief supporting details are also provided to justify these assessments. Further details can be found in the Supplementary Information.
@@
All participants began by completing a brief survey, followed by exposure to their respective experimental conditions, followed by another brief survey and debriefing. Regardless of the condition, all participants saw the same headlines in random order. These stimuli were presented simultaneously with fact-checking information or questions about viewing fact checks (depending on experimental condition) along with questions regarding individual belief and sharing intention. Participants who failed an attention check were excluded from the study. Further details can be found in the Supplementary Information.
Preregistration
Our preregistration (50) included the analysis plan and predicted outcomes related to results presented in the Ineffectiveness of LLM intervention section, excluding the human fact checks condition. Data for this condition was collected later at the request of reviewers. The preregistration also included various exploratory analyses without specific outcome predictions. For all details, we refer the reader to the original preregistration document.
@@ Acknowledgements
The authors would like to acknowledge Lisa Fazio, David Rand, Stephan Lewandowsky, Adam Berinsky, Mark Susmann, and Gordon Pennycook, who kindly shared materials, as well as Byungkyu Lee, for helpful discussions during the early stages of this project. This research was supported in part by Knight Foundation and Volkswagen Foundation.
@@ 1.1 Sampling details
In our final sample, females comprised 53.40% of the sample, males 46.46%, and other genders 0.14%. Age segments were 65+ (20.66%), 55-64 (16.78%), 45-54 (17.74%), 35-44 (16.91%), 25-34 (18.25%), and 18-24 (9.68%). Race percentages were: White (60.17%), Hispanic or Latino/a (17.46%), Black or African American (13.43%), Asian (5.51%), and Other (3.43%). Slightly more than half of the sample (51.92%) had less than a college education, while 48.08% had a college degree. With respect to party identification, 50.72% identified as Democrat or Democrat-leaning, 43.68% as Republican or Republican-leaning, and 5.60% as Independent.
The sampling plan for the control, LLM-optional, and LLM-forced conditions, for both the belief and sharing groups, was preregistered(50) with the goal of obtaining .95 power to detect a small effect size of .1 at the standard .05 error probability with two-by-three-level between-subject manipulations ([Belief vs. Sharing groups] $\times$ [Control, LLM-forced, LLM-optional conditions]). Power analysis by the G*Power(80) software suggested a minimum number of 44 subjects per condition ($N=264$), but we aimed for a larger target sample size of $N=1{,}500$ (250 participants per condition) to increase the precision of our measurements. As noted in the Materials and Methods, data for the human fact check conditions were gathered later at reviewers’ request. Gathering this data required larger sample sizes to meet the minimum spending threshold of our survey partner (Qualtrics) for both the belief and sharing conditions (300 participants per condition).
@@ 1.2 Stimuli curation
The news headlines used as stimuli were selected from a project aimed at comparing misinformation interventions(79). Specifically, 40 headlines were selected from a set of 317 political news stories using a pretest approach(81, 82, 83) to balance the selected headlines in terms of perceived partisanship, impact, familiarity, sensationalism, and the likelihood of being shared and believed.
@@
The claims were presented in a digital format resembling popular social media platforms, commonly known as the “Facebook format”(51), which includes an image, the article headline, and a lede sentence (if present). See the Headlines and fact checks section for all stimuli text.
1.3 LLM fact check generation
@@
A new ChatGPT session was created on the publicly available OpenAI website (chat.openai.com), where the headline text was inserted into a prompt asking, “I saw something today that claimed $<$HEADLINE TEXT$>$. Do you think that this is likely to be true?” The source of an article (e.g., “nytimes.com”) was excluded. If an article’s lede sentence was shown in the stimulus image, it was also included in the prompt, separated by a colon. The prompt for each headline was provided to ChatGPT only once, and the response was saved as a screenshot. All headlines were generated on January 25, 2023, between 12:30–8:00pm Eastern Standard Time. According to the release notes(84), the language model utilized by ChatGPT at that time was a version of GPT-3.5 that has since been updated and is no longer available. See the Headlines and fact checks section for the text of all fact checks as well as the Accuracy of different prompt methods section for further analysis of model accuracy.
1.4 Human fact check generation
Human fact checks were generated to create a uniform structure with clear judgments, as outlined in the Materials and Methods. Fact checks for false headlines were selected from the same misinformation intervention study(79) from which headline stimuli were selected. Since that study did not create fact checks for true headlines, one of the authors manually created these by reading each article to identify accurate and relevant information to support the veracity label. See the Headlines and fact checks section for the text of all fact checks.
1.5 Attrition
Drop out rates varied between 1%–6% across experimental conditions, as reported in Table S1.
Condition Drop-out Attention-check failure
Belief control 1.46% 63.25%
Belief LLM-forced 3.98% 55.06%
Belief LLM-optional 5.27% 54.26%
Belief human fact check 3.33% 67.32%
Sharing control 1.80% 50.27%
Sharing LLM-forced 6.06% n/a
Sharing LLM-optional 3.62% 54.96%
Sharing human fact check 3.19% 60.05%
Table S1: Drop out and attention-check failure rates for each experimental condition.
@@ 1.5 Attrition
We incorporated an attention-check question that involved a headline created by the researchers stating that the color of the sky is yellow. Prior to viewing any headlines, participants were informed about this specific headline and instructed to later answer “Yes” when asked if they believed the headline or were willing to share it, depending on their respective experimental conditions. To minimize the distinction between the attention check and the regular experimental stimuli, this question was formatted in the same manner as all other headlines. This attention check headline was then presented randomly within the 40 stimuli headlines. Participants who answered this question incorrectly were not included in analyses. Table S1 reports on the attention-check failure rates in the different groups. In one group this rate is not available due to a Qualtrics data collection error.
Using $\chi^{2}$ tests, we compared the attrition rates between the control and experimental conditions (LLM-forced, LLM-optional, human fact check). The LLM-forced condition within the sharing group was excluded from this analysis due to the data collection issues mentioned above. This analysis revealed significant differences in attrition between the control and human fact check conditions in the sharing group (Bonferroni adjusted $P<0.001$). Despite matching experimental groups on key demographic characteristics and maintaining identical experimental protocols, these attrition differences may have resulted from different participant recruitment procedures employed by Qualtrics at different times. No other evidence of differential attrition was found.
Table S2 lists the number of participants who were screened out for other reasons prior to being assigned to an experimental group.
Screen-out type Num. of participants
Did not consent 551
Age ($<$ 18 y/o) 58
Non-US resident 37
Would not agree to give their best answers 86
Table S2: Screen-out attrition. These participants were never assigned to an experimental group.
@@ 2.2 Attitude towards AI
Questions were answered with a seven-point Likert scale ranging from “strongly disagree” to “strongly agree.” Items 1 and 3 were reverse coded such that higher values on all items indicated greater trust in artificial intelligence. For our regression analyses (see the Regression analyses section for details), each participant’s ATAI is calculated as the mean value of their responses to this battery.
@@ 2.3 Headline congruence
A headline is considered “congruent” with a participant’s partisan perspective if it is typically considered to be favorable towards the political party that they are affiliated with. Headlines are either pro-Democrat or pro-Republican, based on the pretest described in the main text. Thus, a congruent headline for a Democrat (Republican) would be one that is pro-Democrat (pro-Republican). Conversely, an incongruent headline for a Democrat (Republican) would be one that is pro-Republican (pro-Democrat). In regression analyses, we recode congruent headlines as 0 and incongruent headlines as 1.
We estimated partisanship by asking participants the following question: “Generally speaking do you think of yourself as a Republican, a Democrat, an Independent, or what?” Possible answers were “Democrat,” “Republican,” “Independent,” “No Preference,” “Don’t know,” and “Other” (with a text box to fill if this option is selected). If “Democrat” or “Republican” was not selected as their answer to this question they were then asked, “Do you think of yourself as closer to the Republican or Democratic Party?” Possible answers were “Republican Party,” “Democratic Party,” “Don’t know,” and “Neither.” We consider participants who answered “Democrat” for the first question or “Democratic Party” for the second question as Democrats. We consider as Republicans those who answered “Republican” for the first question or “Republican Party” for the second question. In other words, those who lean towards Democrats (Republicans) were recoded as Democrats (Republicans) in our analysis. All others are considered Independents.